OpenStudy (anonymous):

Solve the system of equations. Show your work. x + 2y = 5 3x + 2y = 17 im still beyond confused... :l

OpenStudy (anonymous):

lol no u will show me ur work :P ok lets use the substitution method, familiar with that one?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

cool can u do that now then? or would u like me to start u off?

OpenStudy (anonymous):

i can probably do it now.. thanks

OpenStudy (anonymous):

What you have to do is just eliminate the "y" just by subtracting those two equations.

OpenStudy (anonymous):

if u want u can show us ur working and we will correct it or show us ur answers and we will confirm them :)

OpenStudy (anonymous):

lol, that's a better idea @HawkCrimson

OpenStudy (anonymous):

the answer i got for the first equation is y = 2.5 + -0.5x

OpenStudy (anonymous):

Im still new at this

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

ok and y = 8.5 + -1.5x is what i got for the second one

OpenStudy (anonymous):

umm u want to have exact answers for y and x

OpenStudy (anonymous):

yep that's right

OpenStudy (anonymous):

like y = 5 and x=3

OpenStudy (anonymous):

so wat u did didnt answer the question.. so u know how to solve for y and x?

OpenStudy (anonymous):

it's still the same, but yeah you should put it as \[\frac{ 5 }{ 2 }\]and all the rest in fractions

OpenStudy (anonymous):

Divide both sides by two. right?

OpenStudy (anonymous):

yes @mirandabell902

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y = 5/2 for the first and y = 5/8 for the second one, right?

OpenStudy (anonymous):

both of them r supposed to have the same y and x value

OpenStudy (anonymous):

ok ill do it with u cause idk wat u were doing before and i dont know how u got those two equations but ill show u how to do it by substitution :)

OpenStudy (anonymous):

first take the 1st equation x + 2y = 5 and make y or x the subject of the formula (it doesnt matter)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wat did u get?

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

calm down, one step ata time and i didnt get that.. just rearrange the formula for me and show me how it looks with x or y being the subject of the formula.

OpenStudy (anonymous):

y=2+5

OpenStudy (anonymous):

x + 2y = 5 x = 5 - 2y or y = (5-x)/2 that is wat i wanted, understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so choose any one lol

OpenStudy (anonymous):

x = 5 - 2y

OpenStudy (anonymous):

The easiest way is just subtract. that way, it cancels 2y in the first equation and 2y in the secnd equation. then, you'll get the x

OpenStudy (anonymous):

yea..nice to know @Lynncake .. as i was saying @mirandabell902 :P

OpenStudy (anonymous):

\[-\left(\begin{matrix}x+2y=1 \\ 2x+2y=1\end{matrix}\right)\]\[-x=0 \rightarrow x=0\]

OpenStudy (anonymous):

or u can interupt me again...sure.

OpenStudy (anonymous):

:D sorry, ok go on

OpenStudy (anonymous):

lol even im confused now D:

OpenStudy (anonymous):

eh just show her the subtraction method lol

OpenStudy (anonymous):

but we should be confusing her with showing her two methods at the same time lol

OpenStudy (anonymous):

yeah you're right, but just finish yours :D @HawkCrimson

OpenStudy (anonymous):

i have to go bye. u do the rest and plus she isnt even replying..